Young's Rule for calculating a child's dosage of medication is
CHILD'S DOSE EQUALS THE AGE OF CHILD DIVIDED BY THE AGE OF CHILD PLUS 12 TIMES ADULT DOSE.

If a 12year old child's dose of a particular medication is 15mg, what is the adult's dose?

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

CHILD'S DOSE EQUALS THE AGE OF CHILD DIVIDED BY THE AGE OF CHILD PLUS 12 TIMES ADULT DOSE.

If a 12year old child's dose of a particular medication is 15mg, what is the adult's dose?
-------------------------
CD = CA/(CA+12S)
15mg = 12/(12 + 12AD)
15mg = 1/(1+AD)
15 + 15AD = 1
15AD = -14
AD = -14/15
-----------------
That doesn't make any sense.
Did you post the correct information?
Cheers,
Stan H.

Answer by solver91311(24713)





Solve for :



Substitute values:



Do the arithmetic. You can do that yourself.

John